My question is how can i find an easier way to tell which problems are associative and commutative. Whenever I see a problem and tells me to name the property I get really confused.
Answer by Theo(2978)

commutative associative distributive
-----
commutative has to do with order.
b+a = a+b
a*b = b*a
if you want to add 5 + 3, you can add 5 + 3 or you can add 3 + 5 and you will get 8 either way.
-----
associative has to do with grouping.
a + b + c = a + (b + c) = (a + b) + c
a * b * c = a * (b * c) = (a * b) * c
if you want to add 5 + 3 + 6 you can add 5 + 3 first to get 8 and then you can add 6 to get 14. if you add 3 + 6 first to get 9 and you add 9 + 5, you still get 14.
-----
distributive has to do with being able to remove a parentheses by distributing the operation amongst the members.
a * (b + c) = a*b + a*c
(a + b) * c = a*c + b*c
if you multiply 5 * (6 + 3) you add the terms in the parentheses first and then multiply to get 5 * 9 = 45. if you remove the parentheses first, you get 5*6 + 5*3 equals 30 + 15 = 45 just as well.
-----
a good website to reference that explains this simply would be:
http://www.onlinemathlearning.com/number-properties.html
-----